OpenStudy (zubhanwc3):

Find the points on the graph of y= sec x, 0<=x<=2pi, where the tangent is parallel to the line 3y-2x=4

OpenStudy (freckles):

have you found y'?

OpenStudy (freckles):

you need to find the slope of 3y-2x=4 and set it equal to the d/dx sec(x) and the solve

OpenStudy (freckles):

that is, \[\frac{d}{dx}\sec(x)=\text{ find slope of } (3y-2x=4)\]

OpenStudy (freckles):

the solve that equation

OpenStudy (zubhanwc3):

alright so y= 2/3x +4/3 y' = 2/3 z= sec x z' = sec x tan x sec x tan x = 2/3 so i should find the value of x so that z' = 2/3?

OpenStudy (zubhanwc3):

in this case it should be pi/6, correct?

OpenStudy (zubhanwc3):

@freckles

OpenStudy (freckles):

yes we are looking for when sec(x)tan(x)=2/3 we are doing that because the slopes need to be the same parallel lines have same slopes (different y-intercept though )

OpenStudy (zubhanwc3):

ah, but do we need to solve for x then?

OpenStudy (freckles):

yes

OpenStudy (zubhanwc3):

or should we rewrite it to 3tanxsecx = 2

OpenStudy (freckles):

you can rewrite it

OpenStudy (zubhanwc3):

o wait, it says find the point, so pi/6 is the answer then, right?

OpenStudy (freckles):

hmm well tan(pi/6)=1/sqrt(3) and sec(pi/6)=2

OpenStudy (freckles):

i'm thinking of turning it into an algebraic equation

OpenStudy (freckles):

let tan(x)=u  sec(x)=sqrt(u^2+1)